OpenStudy (anonymous):

Find the distance between the cities. Assume that Earth is a sphere of radius 4000 miles and that the cities are on the same longitude (one city is due north of the other). (Round your answer to one decimal place.) City Latitude Dallas, Texas 32° 47' 39'' N Omaha, Nebraska 41° 15' 50'' N

OpenStudy (aaronq):

i'm not familiar with these coordinates, but it's essentially finding the arc length.  Since they're on the same longitude (east-west), we consider them to be on the same plane. so arc length=\(\large s=\theta *r=(\theta _1-\theta _2)*r\) you need to convert the degrees given to radians first though e.g. \(\large 32^o*\color{red}{\dfrac{\pi}{180^o}}=0.56~ rad\)

OpenStudy (anonymous):

Thanks.. Can you also tell me how to deal with angles given in this 32° 47' 39'' form.

OpenStudy (aaronq):

No problem. I'm assuming that for this question the latter part (47' 39'') doesn't matter because they are on the same latitude. I've also never used geographical coordinate systems so i have no idea. If you're interested take a look at this http://downloads2.esri.com/ESRIpress/images/205/121250-MapUse-samplechapter.pdf

OpenStudy (anonymous):

Thanks aloott. You're a lifesaver dude :D

OpenStudy (aaronq):

no problem man!